# Insulator phases of Bose–Fermi mixtures induced by intraspecies next-neighbor interactions

### v2

@@ 5 Conclusions
In the present study, we have introduced three incommensurable insulators that emerge after taking into account next-neighbor intraspecies interactions for a Bose–Fermi mixture model of two-color fermions and scalar bosons at the hardcore limit. We used an MPS-based DMRG method to obtain the ground state for given carrier densities and interaction parameters and subsequently constructed plots of the bosonic density against the bosonic chemical potential to determine the presence of the corresponding incompressible phases. We proposed order parameters, as well as constructed multiple density profiles and phase diagrams, varying the next-neighbor interaction between fermions or bosons to characterize the properties of the cited insulators.
@@
The effect of the non-local interactions on the well-known mixed Mott insulator (MMI) [64, 31] and the spin-selective Mott insulator (SSMI) [65, 66, 67, 68] was studied, noting that those insulators tend to disappear or remain stable, depending on whether the value of the fermionic density is greater or less than $\rho_{\rm{F}}=1/2$ and on the kind of species under non-local interactions. For instance, we observed that for $\rho_{\rm{F}}<1/2$ and next-neighbor interactions between fermions, the MMI (SSMI) phase remains stable (tends to disappear) when the non-local coupling grows, while the opposite happens for $\rho_{\rm{F}}>1/2$. The above scenario is exchanged if we consider non-local interaction between bosons instead of fermions.
@@
An immiscible charge density wave (ICDW) phase emerges at $\rho_{\rm{B}}=1/2$, which has a CDW order for both fermions and bosons that are completely out of phase. It was established that this insulator only appears for $1/2\leq\rho_{\rm{F}}\leq 1/(1+I)$, which is associated with the system having just enough fermions to create the characteristic dimerized lattice from the CDW order. The gap from this incompressible phase grows with increasing next-neighbor fermion interaction, since this non-local repulsion provides stability to the CDW structure. The presence of the phase was clarified by employing an order parameter constructed from a product of the traditional CDW order parameter for each species, with a minus sign that accounts for the out-of-phase oscillation. This quantity agrees with the gap opening and gives a stronger argument for the determination of the critical point. Exchanging fermions with bosons for the next-neighbor interaction removes the left border of this phase, allowing it to occur at any fermionic density.
@@
On the other hand, another non-trivial insulator appears for the condition $\rho_{\rm{B}}-\rho_{\rm{F}}=0$, where both species have CDW orders with a characteristic wave vector proportional to the density of both species, which we denote mixed charge density wave (MCDW). We characterize its critical transition using an order parameter analog to the ICDW one, where we change the wave vector to the corresponding one for the MCDW phase, which we corroborate with the bosonic gap calculations. This phase only appears for $\rho_{\rm{B}}+\rho_{\rm{F}}\leq 1$, and for a fermionic density of $\rho_{\rm{F}}=2/5$ it has a unit cell of length five sites that contains two particles of each species. This phase is favored, and its bosonic gap increases with the next-neighbor interaction between fermions or bosons.
@@
A total number of carriers greater than the lattice size $\left(\rho_{\rm{B}}+\rho_{\rm{F}}\geq 1\right)$ and interactions open the possibility of spin-selective states, a fact that was corroborated by the emergence of the spin-selective charge density wave (SSCDW) insulator when turning on the non-local interaction between fermions or bosons. In this state, one kind of fermion is itinerant, while the other couples with the bosons to establish an insulator state that fulfills $\rho_{\rm{B}}-\rho_{\rm{F}}^{\uparrow,(\downarrow)}=0$ and exhibits interleaved CDW profiles with a global pattern. We explicitly showed the spin-selective character of this state by breaking the $\rm{SU}(2)$ symmetry, i.e., under a spin-population imbalance, this state splits into two SSCDW insulators, which are separated by a superfluid phase. Finally, a spin-selective order parameter was proposed and used to identify the superfluid-SSCDW critical point.
@@
The present study unearthed three different incommensurable insulators induced by non-local interactions; however, some aspects need to be explored in the future. For instance, what happens if the hardcore approximation is relaxed? In this case, we anticipate that the new incommensurable insulators will emerge along with the MMI and SSMI states between every two trivial bosonic Mott insulators ($\rho_{\rm{B}}=n$, with $n$ being an integer). This has already been observed for the MMI and SSMI phases, where the relations that they follow change to $\rho_{\rm{B}}+\rho_{\rm{F}}=n$ and $\rho_{\rm{B}}+\rho^{\uparrow,(\downarrow)}_{\rm{F}}=n$, respectively [66]. Either way, we do not discount the appearance of other insulators that could emerge from reducing the hardcore restriction.
@@
This study focused on a particular set of values for the local interspecies and intraspecies interactions, but it is important to think of the effect that changing these parameters can have on these insulators. By setting $U_{\rm{FF}}$, the insulator states are expected to arise from critical values of the Bose–Fermi couplings, which will depend on each phase, as has been shown in previous studies without non-local interactions [64, 65, 66, 67, 68]. For instance, we observed that for $\rho_{\rm{F}}=4/5$, the stability of the SSCDW and ICDW insulators increases for higher values of the local interspecies interaction $U_{\rm{BF}}$, since this parameter modulates the mutual behavior between species that characterize these CDW insulators. On the other hand, a small increase in the local fermionic interaction $U_{\rm{FF}}$ can lead to a charge gap increase in the ICDW phase, due to the localization of the fermions in the CDW order, and a decrease in the SSCDW phase stability because of stronger perturbations caused by the free fermions on the insulator. However, a more thorough study has yet to be done.
@@
The ground state evolution under the simultaneous effect of the next-neighbor interaction between fermions or bosons and/or the inclusion of interspecies non-local couplings was not thoroughly explored. However, some preliminary results suggest that the insulator states unearthed in this paper will emerge. Moreover, we do not discount the possibility of finding new states under these conditions.
@@
We trust that our investigation can inspire future research on the extended Bose-Fermi Hubbard model and specifically the previously-shown incompressible phases, since their incommensurable CDW character presents an opportunity to observe new emerging behavior in mixtures that could give rise to possible applications in industry [95]. Fortunately, the ICDW phase has recently been observed in the excitonic system of GaAs bilayers [34], so an experimental observation of the other insulators would be a crucial input for the research on this topic. The latter could be done in the same condensed matter system where the non-local interactions of the dipolar excitons play a crucial role [33, 34, 35, 36, 37], while cold-atom setups also present a great opportunity. Here the non-local interaction can be fulfilled using polar molecules [96, 97] which could potentially be loaded in an optical lattice, or Rydberg-dressed atoms [98, 99, 100], where proposals of extended Hubbard models have been made. These possible experimental realizations, along with the present proposal of phases with non-local interactions open a pathway to an amalgam of new strongly-correlated insulators in low-dimensional systems.
@@ Appendix A Particle-hole symmetry
In this appendix, we will show that the Hamiltonian (1)-(4) has a particle-hole symmetry at the hardcore limit for the bosons, which allows us to simplify the phase diagrams in Fig. 1. First, we consider the symmetry of the fermions in the system and then we address the case for the bosons.
We introduce the charge conjugation operator $\hat{\mathcal{P}}$, which changes each filled Fock state for a vacuum one and vice-versa
@@
Because of the Pauli exclusion principle, this operator is well-defined for each fermionic site. To generalize the concept for a set of fermionic sites, the charge conjugation operator is defined through its effect on the creation and annihilation operators $\hat{f}_{\sigma,j}^{\dagger}$ and $\hat{f}_{\sigma,j}$, respectively [101]
@@
Now let us show that the fermionic Hamiltonian (3) is invariant under the effect of $\hat{\mathcal{P}}$. The fermionic hopping term directly shows this property
@@
In this case, the interaction exchange is not directly invariant under the transformation. Nevertheless, the extra terms that appear are either constant or in the form of a chemical potential contribution. Since we are working in a number-conserving Hilbert space, they do not change the ground state of the Hamiltonian, as their main effect is to shift the zero in the energy and chemical potential scales. For the fermionic next-neighbor interaction, we recover an analog result
@@
For the bosons, the charge conjugation operator (10) and (11) is ill-defined, since there can be more than two states per site; nevertheless, for the hardcore limit we extend the definition to the bosonic Hilbert space as
@@
where we omit the $(-1)^{j}$ factor due to the bosonic nature of the many-body wave function. This does not affect the result from (14)-(15), since the hardcore bosons have the same on-site anti-commutation relations as the fermions, hence $\hat{\mathcal{P}}\hat{n}^{\rm{B}}_{j}\hat{\mathcal{P}}^{-1}=1-\hat{n}^{\rm{B}}_{j}$.
With (22)-(23), it is clear to show that
@@
So the bosonic Hamiltonian (2) is also invariant under this extended charge conjugation operator.
When we look at the interspecies Hamiltonian (4), it can be seen that both definitions (12)-(13) and (22)-(23) together have the following effect on the interaction term
@@
Hence, we conclude that (1)-(4) is invariant under the charge conjugation transformation up to constant and chemical potential terms. This means that if we find the ground state corresponding to a given set of parameters and densities of particles $\rho_{\rm{B}}$, $\rho_{\rm{F}}^{\uparrow}$ and $\rho_{\rm{F}}^{\downarrow}$, because of the particle-hole symmetry this will also be the ground state for the fillings $1-\rho_{\rm{B}}$, $1-\rho_{\rm{F}}^{\uparrow}$ and $1-\rho_{\rm{F}}^{\downarrow}$, respectively. The ground-state energies will differ because of the extra terms in the transformation of the interactions, but this will only apply a shift in both the energy scale and the chemical potential scale in the $\rho_{\rm{B}}-\mu_{\rm{B}}$ curves used in the present investigation, meaning that the physical properties of the system remain invariant. Because of this, we need only analyze half of the phase diagrams from Fig. 1.

### v3

@@ 5 Conclusions
In the present study, we have introduced three CDW insulators that emerge after taking into account next-neighbor intraspecies interactions for a Bose–Fermi mixture model of two-color fermions and scalar bosons at the hardcore limit. We used an MPS-based DMRG method to obtain the ground state for given carrier densities and interaction parameters and subsequently constructed plots of the bosonic density against the bosonic chemical potential to determine the presence of the corresponding incompressible phases. We proposed order parameters, as well as constructed multiple density profiles and phase diagrams, varying the next-neighbor interaction between fermions or bosons to characterize the properties of the cited insulators.
@@
The effect of the long-ranged interactions on the well-known mixed Mott insulator (MMI) [66, 31] and the spin-selective Mott insulator (SSMI) [67, 68, 69, 70] was studied, noting that those insulators tend to disappear or remain stable, depending on whether the value of the fermionic density is greater or less than $\rho_{\rm{F}}=1/2$ and the kind of species under long-ranged interactions. For instance, we observed that for $\rho_{\rm{F}}<1/2$ and next-neighbor interactions between fermions, the MMI (SSMI) phase remains stable (tends to disappear) when the long-ranged coupling grows, while the opposite happens for $\rho_{\rm{F}}>1/2$. The above scenario is exchanged if we consider long-ranged interaction between bosons instead of fermions.
@@
An immiscible charge density wave (ICDW) phase emerges at $\rho_{\rm{B}}=1/2$, which has a CDW order for both fermions and bosons that are completely out of phase. It was established that this insulator only appears for $1/2\leq\rho_{\rm{F}}\leq 1/(1+I)$, which is associated with the system having just enough fermions to create the characteristic dimerized lattice from the CDW order. The gap from this incompressible phase grows with increasing next-neighbor fermion interaction since this long-ranged repulsion provides stability to the CDW structure. The presence of the phase was clarified by employing an order parameter constructed from a product of the traditional CDW order parameter for each species, with a minus sign that accounts for the out-of-phase oscillation. This quantity agrees with the gap opening and gives a stronger argument for the determination of the critical point. Exchanging fermions with bosons for the next-neighbor interaction removes the left border of this phase, allowing this phase to occur at any fermionic density.
@@
On the other hand, another non-trivial insulator appears for the condition $\rho_{\rm{B}}-\rho_{\rm{F}}=0$, where both species have CDW orders with a characteristic wave vector proportional to the density of both species, which we denote as mixed charge density wave (MCDW). We characterize its critical transition using an order parameter analog to the ICDW one where we change the wave vector to the corresponding one for the MCDW phase, which we corroborate with the bosonic gap calculations. This phase only appears for $\rho_{\rm{B}}+\rho_{\rm{F}}\leq 1$ and for a fermionic density of $\rho_{\rm{F}}=2/5$ it has a unit cell of length five sites that contains two particles of each species. This phase is favored, and its bosonic gap increases with the next-neighbor interaction between fermions or bosons.
@@
A total number of carriers greater than the lattice size $\left(\rho_{\rm{B}}+\rho_{\rm{F}}\geq 1\right)$ and interactions open the possibility to spin-selective states, a fact that was corroborated by the emergence of the spin-selective charge density wave (SSCDW) insulator when turning on the long-ranged interaction between fermions or bosons. In this state, one kind of fermion is itinerant, while the other couples with the bosons to establish an insulator state that fulfills $\rho_{\rm{B}}-\rho_{\rm{F}}^{\uparrow,(\downarrow)}=0$ and exhibits interleaved CDW profiles with a global pattern. We showed explicitly the spin-selective character of this state by breaking the $\rm{SU}(2)$ symmetry, i.e., under a spin-population imbalance, this state splits into two SSCDW insulators, which are separated by a superfluid phase. Finally, a spin-selective order parameter was proposed and used to identify the superfluid-SSCDW critical point.
@@
The present study unearthed three different CDW insulators induced by long-ranged interactions, however, some aspects need to be explored in the future. For instance, what happens if the hardcore approximation is relaxed? In this case, we anticipate that the new CDW insulators will emerge along with the MMI and SSMI states between every two trivial bosonic Mott insulators ($\rho_{\rm{B}}=n$, with $n$ being an integer). This has already been observed for the MMI and SSMI phases, where the relations that they follow change to $\rho_{\rm{B}}+\rho_{\rm{F}}=n$ and $\rho_{\rm{B}}+\rho^{\uparrow,(\downarrow)}_{\rm{F}}=n$, respectively [68]. Either way, we do not discount the appearance of other insulators that could emerge from reducing the hardcore restriction.
@@
This study is focused on a particular set of values for the local interspecies and intraspecies interactions, but it is important to think of the effect that changing these parameters can have on these insulators. By setting $U_{\rm{FF}}$, the insulator states are expected to arise from critical values of the Bose–Fermi couplings, which will depend on each phase, as has been shown in previous studies without long-ranged interactions [66, 67, 68, 69, 70]. For instance, we observed that for $\rho_{\rm{F}}=4/5$, the stability of the SSCDW and ICDW insulators increases for higher values of the local interspecies interaction $U_{\rm{BF}}$, since this parameter modulates the mutual behavior between species that characterize these CDW insulators. On the other hand, a small increase in the local fermionic interaction $U_{\rm{FF}}$ can lead to a charge gap increase in the ICDW phase, due to the localization of the fermions in the CDW order, and a decrease in the SSCDW phase stability because of stronger perturbations caused by the free fermions on the insulator. However, a more thorough study has yet to be done.
@@
The ground state evolution under the simultaneous effect of the next-neighbor interaction between fermions or bosons and/or the inclusion of interspecies long-ranged couplings was not widely explored. However, some preliminary results suggest that the insulator states unearthed in this paper will emerge. Moreover, we do not disregard the possibility of finding new states under these conditions. Nevertheless, this falls outside the scope of this work.
@@
We trust that our investigation can inspire future research on the extended Bose–Fermi Hubbard model and specifically the previously shown incompressible phases since their CDW character presents an opportunity to observe new emerging behavior in mixtures that could give rise to possible applications in industry [100]. Fortunately, the ICDW phase has been observed in the excitonic system of GaAs bilayers recently [34], then an experimental observation of the other insulators would be a crucial input for the research on this topic. The latter could be done in the same condensed matter system where the long-ranged interactions of the dipolar excitons play a crucial role [33, 34, 35, 36, 37], while cold-atom setups also present a great opportunity, here the long-ranged interaction can be fulfilled using polar molecules [101, 102] which could be potentially loaded in an optical lattice, or Rydberg-dressed atoms [103, 104, 105] where proposals of extended Hubbard models have been made. These possible experimental realizations along with the present proposal of phases with long-ranged interactions open the pathway to an amalgam of new strongly-correlated insulators in low-dimensional systems.
@@ Appendix A Particle-hole symmetry
In this appendix, we will show that the Hamiltonian (1)-(4) has a particle-hole symmetry in the hardcore limit for the bosons which allows us to simplify the phase diagrams in Fig. 1. First, we consider the symmetry of the fermions in the system and then we address the case for the bosons.
We introduce the charge conjugation operator $\hat{\mathcal{P}}$ which changes each filled Fock state for a vacuum one and vice-versa
@@
Because of the Pauli exclusion principle, this operator is well-defined for each fermionic site. To generalize the concept for a set of fermionic sites the charge conjugation operator is defined through its effect on the creation and annihilation operators $\hat{f}_{\sigma,j}^{\dagger}$ and $\hat{f}_{\sigma,j}$, respectively [106]
@@
Now let us show that the fermionic Hamiltonian (3) is invariant under the effect of $\hat{\mathcal{P}}$. The fermionic hopping term shows directly this property
@@
In this case, the interaction exchange is not directly invariant under the transformation. Nevertheless, the extra terms that appear are either constant or of the form of a chemical potential contribution, since we are working in a number-conserving Hilbert space these do not change the ground state of the Hamiltonian as their main effect is to shift the zero in the energy and chemical potential scales. For the fermionic next-neighbor interaction, we recover an analog result
@@
For the bosons the charge conjugation operator (12) and (13) is ill-defined since there can be more than two states per site, nevertheless for the hardcore limit we extend the definition to the bosonic Hilbert space as
@@
where we omit the $(-1)^{j}$ factor due to the bosonic nature of the many-body wave function. This does not affect the result from (16)-(17) since the hardcore bosons have the same on-site anti-commutation relations as the fermions, hence $\hat{\mathcal{P}}\hat{n}^{\rm{B}}_{j}\hat{\mathcal{P}}^{-1}=1-\hat{n}^{\rm{B}}_{j}$.
With (24)-(25) it is direct to show that
@@
Then, the bosonic Hamiltonian (2) is also invariant under this extended charge conjugation operator.
When we look at the interspecies Hamiltonian (4) it is noted that both definitions (14)-(15) and (24)-(25) together have the following effect on the interaction term
@@
Hence, we conclude that (1)-(4) is invariant under the charge conjugation transformation up to constant and chemical potential terms. This means that if we find the corresponding ground state to a given set of parameters and densities of particles $\rho_{\rm{B}}$, $\rho_{\rm{F}}^{\uparrow}$ and $\rho_{\rm{F}}^{\downarrow}$ because of the particle-hole symmetry this will also be the ground state for the fillings $1-\rho_{\rm{B}}$, $1-\rho_{\rm{F}}^{\uparrow}$ and $1-\rho_{\rm{F}}^{\downarrow}$, respectively. The ground-state energies will differ because of the extra terms in the transformation of the interactions, but this will only apply a shift in both the energy scale and the chemical potential scale in the $\rho_{\rm{B}}-\mu_{\rm{B}}$ curves used along the present work, meaning that the physical properties of the system remain invariant. Because of this we only need to analyze half of the phase diagrams from Fig. 1.
Appendix B Order of the phase transitions
In the main text, three different CDW insulators are characterized, and we suggest that future work should consider the study of the associated superfluid-insulator transition. Here, we provide a first insight into this topic by analyzing the system’s ground-state energy as we increase the model’s long-range interactions.
For each of the three fermionic densities from Sec. 3.13.23.3 we plot in Fig. 12 the ground state energy $E$ scaled by the system size and extrapolated to the thermodynamic limit as a function of $V_{\rm FF}\in[0,2]$ for each corresponding highlighted CDW insulator, which are the ICDW ($\rho_{\rm F}=2/3,\rho_{\rm B}=1/2$), MCDW ($\rho_{\rm F}=2/5,\rho_{\rm B}=2/5$) and the SSCDW ($\rho_{\rm F}=4/5,\rho_{\rm B}=2/5$), respectively. At $V_{\rm FF}=0$ the systems start in a superfluid phase, and as we increase $V_{\rm FF}$ the superfluid-insulator transition occurs at a critical interaction showed as a vertical line of the corresponding color and line style. In all cases shown, the energy curve is smooth across the transition, from which we discard a first-order transition where a discontinuity in the energy’s first derivative would generate a nudge in the curves. This agrees with previous literature on CDW phases which affirm its transition to be of second order [107]. Even then, a more meticulous study should be developed in the future for the present mixture system to characterize the nature of the phase transition.
Figure 12: Ground state energy $E$ scaled by the system size in the thermodynamic limit as a function of $V_{\rm FF}$. Three density combinations are shown, each associated with the emergence of a CDW insulator, corresponding to $\rho_{\rm F}=2/3$, $\rho_{\rm B}=1/2$ (purple, ICDW), $\rho_{\rm F}=2/5$, $\rho_{\rm B}=2/5$ (orange, MCDW) and $\rho_{\rm F}=4/5$, $\rho_{\rm B}=2/5$ (cyan, SSCDW), with onsite interactions $U_{\rm FF}=4$ and $U_{\rm BF}=6$. Vertical lines with the corresponding color and line style show the critical point of the transition for each insulating state, obtained in Sec. 3.13.23.3.